OpenStudy (anonymous):

(x^2 + 6x + 9) / (5x +15

Parth (parthkohli):

I see that \(x^2 + 6x + 9\) is a perfect square.

OpenStudy (anonymous):

can you hep me solve

Parth (parthkohli):

\(x^2 + 6x + 9 = (x + 3)^2\) \(5x + 15 = 5(x + 3)\) Something would get cancelled here ;P

OpenStudy (anonymous):

\[x^2+6x+9= (x+3)(x+3)\]

OpenStudy (anonymous):

Well the things that similar will cancel and by cancel we mean it turns into one. Remember anything divided by itself is equal to one!

OpenStudy (anonymous):

so you cancle out (x + 3) and it would be x^2 + 6x + 9 / (x + 3?